# bibCyclic.bib

### v2

@@ 2.6 Wreath Macdonald polynomials
Wreath Macdonald polynomials were originally defined by Haiman in [Haiman1]. Their existence was proven using geometric techniques in [BFwreath]. We review their definition, following the expositions given in [OrrShim] and [Wen].
@@ 2.8 Cauchy identities
The following evaluation formula was recently proven by Wen and Romero.
Proposition 2.7 ([josh]).
Let $10\leq m\leq l-$ 1. If $\text{core}(\lambda)=\emptyset$, then
@@ 3.1 Cyclic quiver varieties
Let $Q$ be the $\hat{A}_{l-1}$ quiver, which has vertex set $I=\{0,1,\ldots,l-1\}$ and edges $i\to i+1$ for $10\leq i\leq l-$ 1. Here and elsewhere, we use the convention that indices corresponding to vertices of $Q$ are taken modulo $l$.
@@ 3.2 Hilbert schemes of points
The Hilbert scheme of $m$ points in $\mathbb{C}^{2}$, written $\text{Hilb}^{m}(\mathbb{C}^{2})$, is the moduli space of ideals $J\subset\mathbb{C}[x,y]$ such that $\dim_{\mathbb{C}}(\mathbb{C}[x,y]/J)=m$. It can be constructed as a Nakajima variety for the Jordan quiver with $\mathsf{v}=(m)$ and $\mathsf{w}=(1)$..
@@ 3.3 Torus action and fixed points
Let $\mathsf{A}=(\mathbb{C}^{\times})^{|\mathsf{w}|}$. There is an action of the torus $\mathsf{T}=\mathsf{A}\times(\mathbb{C}^{\times})^{2}$ on $\mathcal{M}(\mathsf{v},\mathsf{w})$. The $\mathsf{A}$ component acts on $\mathcal{M}$ by its natural diagonal action on $W_{i}$. The second component of $\mathsf{T}$ acts by
@@
Next we describe the torus fixed points of $\mathcal{M}(\mathsf{v},\mathsf{w})$, which relies on the “tensor product” of quiver varieties, see [Naktens]. Choose a decomposition $\mathsf{w}=\mathsf{w}^{\prime}+\mathsf{w}^{\prime\prime}$ corresponding to a choice of splitting $W_{i}=W_{i}^{\prime}\oplus W_{i}^{\prime\prime}$ and let $\mathsf{A}^{\prime}\subset\mathsf{A}$ act trivially on $W_{i}^{\prime\prime}$ and with weight $1$ on $W_{i}^{\prime}$. Letting $a$ be the coordinate on $\mathsf{A}^{\prime}$, we abbreviate this setup by writing $\mathsf{w}=a\mathsf{w}^{\prime}+\mathsf{w}^{\prime\prime}$. Then
@@
Iterating this decomposition, one sees that $\mathsf{T}$-fixed points on $\mathcal{M}(\mathsf{v},\mathsf{w})$ are given by products of (×$(\mathbb{C}^{\times})^{2}$-fixed points on $\mathcal{M}(\mathsf{v}^{\prime},\mathsf{w}^{\prime})$ where $|\mathsf{w}^{\prime}|=1$.
Due to the rotational symmetry of the quiver, it is sufficient to describe $\mathcal{M}(\mathsf{v},\delta_{0})^{(\mathbb{C}^{\times})^{2}}$ where $\delta_{0}=(1,0,\ldots,0)$. It is known that (×$(\mathbb{C}^{\times})^{2}$-fixed points on $\text{Hilb}^{m}(\mathbb{C}^{2})$ are in natural bijection with partitions $\lambda$ of $m$. Since the embedding $\mathcal{M}(\mathsf{v},\delta_{0})\subset\text{Hilb}^{|\mathsf{v}|}(\mathbb{C}^{2})$ is (×$(\mathbb{C}^{\times})^{2}$ equivariant, the (×$(\mathbb{C}^{\times})^{2}$-fixed points of $\mathcal{M}(\mathsf{v},\delta_{0})$ are given by partitions $\lambda$ of $|\mathsf{v}|$ such that the number of boxes of content $i$ modulo $l$ is equal to $\mathsf{v}_{i}$ for all $i$, see also [combqv].
@@
$\mathcal{M}(\mathsf{v},\mathsf{w})^{\mathsf{T}}$ is in bijection with $|\mathsf{w}|$-tuples of partitions $(\lambda^{(i,j)})_{\begin{subarray}{c}0\leq i\leq l-1\\ 1\leq j\leq\mathsf{w}_{i}\end{subarray}}$ such that the total number of boxes with content [^3] The contents of the boxes in $\lambda^{(i,j)}$ are shifted by $i$ from the usual content function of section 2.1. $i$ modulo $l$ is equal to $\mathsf{v}_{i}$.
@@ 3.5 Connection to symmetric functions
([BFwreath, losev]) There is an isomorphism of graded vector spaces over $\mathbb{Q}(t_{1},t_{2})$
@@
such that $[I_{\lambda}]\mapsto H_{\lambda}\otimes e^{\text{core}_{l}(\lambda)}$ where $[I_{\lambda}]$ is the $K$-theory class of the skyscraper sheaf at the fixed point $\lambda\in\bigsqcup_{\mathsf{v}}\mathcal{M}(\mathsf{v},\delta_{0})^{\mathsf{T}}$. Here $\mathbb{Q}[Q]$ is the group algebra of $Q$, where $Q$ is the finite type $A_{l-1}$ root lattice and $\text{core}_{l}(\lambda)$ is identified with an element of $Q$ by Proposition 2.3.
@@
[…]
@@ 4 Quasimap counts
In this section, we will review the main objects appearing in the $K$-theoretic enumerative geometry of Nakajima varieties. Our main references for this section are [qm, Ok17, OkSm]. Since we will not contributed anything new here, we will pass quickly over the details.
@@ 4.1 Moduli of quasimaps
As in previous sections, we fix $l\geq 0$ and consider quiver varieties $\mathcal{M}:=\mathcal{M}(\mathsf{v},\mathsf{w})$ for the type $\hat{A}_{l-1}$ quiver. For $d\in\mathbb{Z}^{l}$, let $\text{QM}^{d}:=\text{QM}^{d}(\mathbb{P}^{1}\to\mathcal{M})$ denote the moduli space of stable degree $d$ quasimaps from $\mathbb{P}^{1}$ to $\mathcal{M}$, see [Ok17] section 4. In particular, a quasimap from $\mathbb{P}^{1}$ to $\mathcal{M}$ is a map from $\mathbb{P}^{1}$ to the quotient stack $\mathcal{M}^{\text{st}}:=[\mu^{-1}(0)/G_{\mathsf{v}}]$.
@@
There is a subset of the moduli space that we will primarily be interested in: the open subset of quasimaps of degree $d$ nonsingular at 00. By definition, this is $\text{QM}^{d}_{\text{ns }0}=\{f\in\text{QM}^{d}\,\mid\,f(0)\in\mathcal{M}\subset\mathcal{M}^{\text{st}}\}$. By definition, there are $G$-equivariant evaluation maps
@@
By [Ok17] sections 6 and 7, there is a space $\text{QM}^{d}_{\text{rel }0}$ and a commutative diagram
@@
It is known that $K_{G}(\mathcal{M}^{\text{st}})=K_{G\times G_{\mathsf{v}}}(pt)$, which is generated over $K_{G}(pt)$ by symmetric Laurent polynomials in the Chern roots of $\mathcal{V}_{i}$ for $10\leq i\leq l-$ 1. We denote elements of $K_{G\times G_{\mathsf{v}}}(pt)$ by $\tau$. Abusing notation, we denote their restriction to $K_{G}(\mathcal{M})$ by the same symbol.
Moduli spaces of quasimaps come equipped with a perfect obstruction theory, and hence natural virtual classes. We denote by $\widehat{\mathcal{O}}_{\text{vir}}^{d}$ the symmetrized virtual structure sheaf on a moduli space of quasimaps of degree $d$, see [Ok17] section 6. The symmetrization requires a choice of a polarization, which we take to be the virtual bundle
@@ 4.2 Vertex
(Section 7 of [Ok17]) The bare vertex with descendant $\tau\in K_{G\times G_{\mathsf{v}}}(pt)$ is the generating function
@@
In the previous definition, the pushforward $\text{ev}_{0,*}$ is defined by $\mathbb{C}^{\times}_{q}$-localization using diagram (6). In the above definition we use $z^{d}=z_{0}^{d_{0}}\cdots z_{l-1}^{d_{l-1}}$ to keep track of degrees of quasimaps in what are usually called Kähler parameters. The collection of $d\in\mathbb{Z}^{l}$ such that $\text{QM}^{d}\neq\emptyset$ forms a cone the notation $[[z]]$ should be understood as representing a completion of the semigroup algebra of this cone.
For quiver varieties with finitely many torus fixed points and stability condition $\theta=\pm(1,1,\ldots,1)$ (which includes all varieties studied in this paper), bare vertex function can be computed by equivariant localization. For examples of computations in this regard, see [D5Vertex, dinksmir2, liu, dinksmir].
@@ 4.3 Capped vertex
(Section 7 of [Ok17]) The capped vertex with descendant $\tau\in K_{G\times G_{\mathsf{v}}}(pt)$ is the generating function
@@
A remarkable rigidity theorem proven by Okounkov in [Ok17] is that, for sufficiently large framing, the capped vertex function obeys a property known as large framing vanishing.
@@
(Theorem 7.5.23 in [Ok17]) Let $\tau\in K_{G\times G_{\mathsf{v}}}(pt)$. There exists some framing vector $\mathsf{w}(\tau)$ such that for any $\mathsf{w}\geq\mathsf{w}(\tau)$ [^4] Here we mean $\mathsf{w}\geq\mathsf{w}^{\prime}\iff\mathsf{w}_{i}\geq\mathsf{w}^{\prime}_{i}$ for all $i$. the capped vertex with descendant $\tau$ for $\mathcal{M}(\mathsf{v},\mathsf{w})$ is purely classical:
@@ 4.4 Capping operator
There are also quasimap spaces obtained by imposing a nonsingular condition at $\infty$ and a relative condition at 00, denoted by $\text{QM}_{\begin{subarray}{c}\text{rel }0,\\ \text{ns }\infty\end{subarray}}^{d}$.
@@
(Section 8 of [Ok17]) The capping operator is the generating function
@@
In [OkSm], Okounkov and Smirnov showed that the operator $\Psi(z)$ is the fundamental solution of the quantum difference equation. This equation is described by operators in a certain Hopf algebra acting on the equivariant $K$-theory of $\mathcal{M}$.
4.5 The descendants $\tau_{m}$
The following special class of descendants are our main interest.
Definition 4.6.
@@ 5 Large framing vanishing
The proof is the content of the next two subsections.
@@ 5.1 Torus fixed quasimaps and localization
In [Ok17], Okounkov proves that for an arbitrary descendant, the capped vertex will be purely classical for sufficiently large framing. By the logic of the proof in [Ok17] section 7.5, $\hat{V}^{\left(\tau\right)}_{p}=\tau|_{p}$ whenever $\lim_{q\to\infty}V^{\left(\tau\right)}_{p}=\tau|_{p}$.
@@
By (7.5.14) of [Ok17], the $q\to\infty$ limit of the contribution of $f$ to $V^{\left(\tau\right)}$ behaves like
@@
So the $q\to\infty$ limit of the (7) will be 00 if
@@
As discussed in [Ok17] section 7.5, the dependence of the vertex on the polarization is through a simple shift. For the remainder of this section, we take the polarization to be

### v3

@@ 2.6 Wreath Macdonald polynomials
Wreath Macdonald polynomials were originally defined by Haiman in [25]. Their existence was proved using geometric techniques in [5]. We review their definition, following the expositions given in [49] and [66].
@@ 2.8 Cauchy identities
The following evaluation formula was recently proved by Wen and Romero.
Proposition 2.7 ([64]).
Let $0\leq m\leq l-1$. If $\text{core}(\lambda)=\emptyset$, then
@@ 3.1 Cyclic quiver varieties
Let $Q$ be the $\hat{A}_{l-1}$ quiver, which has vertex set $I=\{0,1,\ldots,l-1\}$ and edges $i\to i+1$ for $0\leq i\leq l-1$. Here and elsewhere, we use the convention that indices corresponding to vertices of $Q$ are taken modulo $l$.
@@ 3.2 Hilbert schemes of points
The Hilbert scheme of $m$ points in $\mathbb{C}^{2}$, written $\text{Hilb}^{m}(\mathbb{C}^{2})$, is the moduli space of ideals $J\subset\mathbb{C}[x,y]$ such that $\dim_{\mathbb{C}}(\mathbb{C}[x,y]/J)=m$. It can be constructed as a Nakajima variety for the Jordan quiver with $\mathsf{v}=(m)$ and $\mathsf{w}=(1)$.
@@ 3.3 Torus action and fixed points
Let $\mathsf{A}=(\mathbb{C}^{\times})^{|\mathsf{w}|}$. There is an action of the torus $\mathsf{T}=\mathsf{A}\times(\mathbb{C}^{\times})^{2}$ on $\mathcal{M}(\mathsf{v},\mathsf{w})$. The action of the $\mathsf{A}$ factor on $\mathcal{M}$ is induced by the action of diagonal matrices on $W_{i}$. The second component of $\mathsf{T}$ acts by
@@
Next we describe the torus fixed points of $\mathcal{M}(\mathsf{v},\mathsf{w})$, which relies on the “tensor product” of quiver varieties, see [42]. Choose a decomposition $\mathsf{w}=\mathsf{w}^{\prime}+\mathsf{w}^{\prime\prime}$ corresponding to a choice of splitting $W_{i}=W_{i}^{\prime}\oplus W_{i}^{\prime\prime}$ and let $\mathsf{A}^{\prime}\subset\mathsf{A}$ act trivially on $W_{i}^{\prime\prime}$ and with weight $1$ on $W_{i}^{\prime}$. Letting $a$ be the coordinate on $\mathsf{A}^{\prime}$, we abbreviate this setup by writing $\mathsf{w}=a\mathsf{w}^{\prime}+\mathsf{w}^{\prime\prime}$. Then
@@
Iterating this decomposition, one sees that $\mathsf{T}$-fixed points on $\mathcal{M}(\mathsf{v},\mathsf{w})$ are given by products of $(\mathbb{C}^{\times})^{2}$-fixed points on $\mathcal{M}(\mathsf{v}^{\prime},\mathsf{w}^{\prime})$ where $|\mathsf{w}^{\prime}|=1$.
Due to the rotational symmetry of the quiver, it is sufficient to describe $\mathcal{M}(\mathsf{v},\delta_{0})^{(\mathbb{C}^{\times})^{2}}$ where $\delta_{0}=(1,0,\ldots,0)$. It is known that $(\mathbb{C}^{\times})^{2}$-fixed points on $\text{Hilb}^{m}(\mathbb{C}^{2})$ are in natural bijection with partitions $\lambda$ of $m$. Since the embedding $\mathcal{M}(\mathsf{v},\delta_{0})\subset\text{Hilb}^{|\mathsf{v}|}(\mathbb{C}^{2})$ is $(\mathbb{C}^{\times})^{2}$ equivariant, the $(\mathbb{C}^{\times})^{2}$-fixed points of $\mathcal{M}(\mathsf{v},\delta_{0})$ are given by partitions $\lambda$ of $|\mathsf{v}|$ such that the number of boxes of content $i$ modulo $l$ is equal to $\mathsf{v}_{i}$ for all $i$, see also [21].
For general $\mathsf{w}$, $\mathcal{M}(\mathsf{v},\mathsf{w})^{\mathsf{T}}$ is thus in bijection with certain $|\mathsf{w}|$-tuples of partitions. We denote such a tuple by $(\lambda^{(i,j)})_{\begin{subarray}{c}0\leq i\leq l-1\\ 1\leq j\leq\mathsf{w}_{i}\end{subarray}}$. Whenever we discuss tuples of partitions of this form labeling a torus fixed point, we will use the convention the content of a box $\square\in\lambda^{(i,j)}$ is equal to $c(\square)+i$, where $c$ is the usual content function from section 2.1.
@@
$\mathcal{M}(\mathsf{v},\mathsf{w})^{\mathsf{T}}$ is in bijection with $|\mathsf{w}|$-tuples of partitions $(\lambda^{(i,j)})_{\begin{subarray}{c}0\leq i\leq l-1\\ 1\leq j\leq\mathsf{w}_{i}\end{subarray}}$ such that the total number of boxes with content $i$ modulo $l$ is equal to $\mathsf{v}_{i}$.
@@ 3.4 Tautological bundles
The following tautological classes are our main interest.
Definition 3.4.
@@ 3.5 Connection to symmetric functions
([5, 33]) There is an isomorphism of graded vector spaces over $\mathbb{Q}(t_{1},t_{2})$
@@
such that $I_{\lambda}]\mapsto H_{\lambda}\otimes e^{\text{core}_{l}(\lambda)}$ where $I_{\lambda}]$ is the $K$-theory class of the skyscraper sheaf at the fixed point $\lambda\in\bigsqcup_{\mathsf{v}}\mathcal{M}(\mathsf{v},\delta_{0})^{\mathsf{T}}$. Here $\mathbb{Q}[Q]$ is the group algebra of $Q$, where $Q$ is the finite type $A_{l-1}$ root lattice and $\text{core}_{l}(\lambda)$ is identified with an element of $Q$ by Proposition 2.3.
@@
in localized equivariant $K$-theory, the left hand side of Lemma 2.8 is simply $\hat{V}^{\left(\tau_{-m}\right)}|_{z_{i}=0,\,0\leq i\leq l-1}$.
@@ 4 Quasimap counts
In this section, we will review the main objects appearing in the $K$-theoretic enumerative geometry of Nakajima varieties. Our main references for this section are [9, 46, 48]. Since we will not contributed anything new here, we will pass quickly over the details.
@@ 4.1 Moduli of quasimaps
As in previous sections, we fix $l\geq 0$ and consider quiver varieties $\mathcal{M}:=\mathcal{M}(\mathsf{v},\mathsf{w})$ for the type $\hat{A}_{l-1}$ quiver. For $d\in\mathbb{Z}^{l}$, let $\text{QM}^{d}:=\text{QM}^{d}(\mathbb{P}^{1}\to\mathcal{M})$ denote the moduli space of stable degree $d$ quasimaps from $\mathbb{P}^{1}$ to $\mathcal{M}$, see [46] section 4. In particular, a quasimap from $\mathbb{P}^{1}$ to $\mathcal{M}$ is a map from $\mathbb{P}^{1}$ to the quotient stack $\mathcal{M}^{\text{st}}:=[\mu^{-1}(0)/G_{\mathsf{v}}]$.
@@
There is a subset of the moduli space that we will primarily be interested in: the open subset of quasimaps of degree $d$ nonsingular at 0. By definition, this is $\text{QM}^{d}_{\text{ns }0}=\{f\in\text{QM}^{d}\,\mid\,f(0)\in\mathcal{M}\subset\mathcal{M}^{\text{st}}\}$. By definition, there are $G$-equivariant evaluation maps
@@
By [46] sections 6 and 7, there is a space $\text{QM}^{d}_{\text{rel }0}$ and a commutative diagram
@@
It is known that $K_{G}(\mathcal{M}^{\text{st}})=K_{G\times G_{\mathsf{v}}}(pt)$, which is generated over $K_{G}(pt)$ by symmetric Laurent polynomials in the Chern roots of $\mathcal{V}_{i}$ for $0\leq i\leq l-1$. We denote elements of $K_{G\times G_{\mathsf{v}}}(pt)$ by $\tau$. Abusing notation, we denote their restriction to $K_{G}(\mathcal{M})$ by the same symbol.
Moduli spaces of quasimaps come equipped with a perfect obstruction theory, and hence natural virtual classes. We denote by $\widehat{\mathcal{O}}_{\text{vir}}^{d}$ the symmetrized virtual structure sheaf on a moduli space of quasimaps of degree $d$, see [46] section 6. The symmetrization requires a choice of a polarization, which we take to be the virtual bundle
@@ 4.2 Vertex
(Section 7 of [46]) The bare vertex with descendant $\tau\in K_{G\times G_{\mathsf{v}}}(pt)$ is the generating function
@@
In the previous definition, the pushforward $\text{ev}_{0,*}$ is defined by $\mathbb{C}^{\times}_{q}$-localization using diagram (8). In the above definition we use $z^{d}=z_{0}^{d_{0}}\cdots z_{l-1}^{d_{l-1}}$ to keep track of degrees of quasimaps in what are usually called Kähler parameters. The collection of $d\in\mathbb{Z}^{l}$ such that $\text{QM}^{d}\neq\emptyset$ forms a cone; the notation $[[z]]$ should be understood as representing a completion of the semigroup algebra of this cone.
For quiver varieties with finitely many torus fixed points and stability condition $\theta=\pm(1,1,\ldots,1)$ (which includes all varieties studied in this paper), bare vertex function can be computed by equivariant localization. For examples of computations in this regard, see [13, 18, 32, 60].
@@ 4.3 Capped vertex
(Section 7 of [46]) The capped vertex with descendant $\tau\in K_{G\times G_{\mathsf{v}}}(pt)$ is the generating function
@@
A remarkable rigidity theorem proved by Okounkov in [46] is that, for sufficiently large framing, the capped vertex function obeys a property known as large framing vanishing.
@@
(Theorem 7.5.23 in [46]) Let $\tau\in K_{G\times G_{\mathsf{v}}}(pt)$. There exists some framing vector $\mathsf{w}(\tau)$ such that for any $\mathsf{w}\geq\mathsf{w}(\tau)$ [^4] Here we mean $\mathsf{w}\geq\mathsf{w}^{\prime}\iff\mathsf{w}_{i}\geq\mathsf{w}^{\prime}_{i}$ for all $i$. the capped vertex with descendant $\tau$ for $\mathcal{M}(\mathsf{v},\mathsf{w})$ is purely classical:
@@ 4.4 Capping operator
There are also quasimap spaces obtained by imposing a nonsingular condition at $\infty$ and a relative condition at 0, denoted by $\text{QM}_{\begin{subarray}{c}\text{rel }0,\\ \text{ns }\infty\end{subarray}}^{d}$.
@@
(Section 8 of [46]) The capping operator is the generating function
@@
In [48], Okounkov and Smirnov showed that the operator $\Psi(z)$ is the fundamental solution of the quantum difference equation. This equation is described by operators in a certain Hopf algebra acting on the equivariant $K$-theory of $\mathcal{M}$.
@@ 5 Large framing vanishing
We will prove this theorem at the end of section 5.3 after setting up the necessary prerequisites in the next three sections.
@@ 5.1 Torus fixed quasimaps and localization
In [46], Okounkov proves that for an arbitrary descendant, the capped vertex will be purely classical for sufficiently large framing. By the logic of the proof in [46] section 7.5, $\hat{V}^{\left(\tau\right)}_{p}=\tau|_{p}$ whenever $\lim_{q\to\infty}V^{\left(\tau\right)}_{p}=\tau|_{p}$.
@@
By (7.5.14) of [46], the $q\to\infty$ limit of the contribution of $f$ to the localization formula for $V^{\left(\tau\right)}$ behaves like
@@
So the $q\to\infty$ limit of the (9) will be 0 if
@@
By localization, it suffice to prove this for $f\in(\text{QM}_{\text{ns }0})^{\mathsf{T}\times\mathbb{C}^{\times}_{q}}$, i.e. for quasimaps fixed additionally by $\mathsf{T}$.
As discussed in [46] section 7.5, the dependence of the vertex on the polarization is through a simple shift. For the remainder of this section, we take the polarization to be
@@
5.2 Torus fixed quasimaps and reverse plane partitions
Let $f\in\text{QM}_{\text{ns }0}\left(\mathbb{P}^{1}\to\mathcal{M}(\mathsf{v},\mathsf{w})\right)^{\mathsf{T}\times\mathbb{C}^{*}_{q}}$ be a quasimap such that $f(0)=x$. Let us recall the combinatorial description of such $f$ in terms of reverse plane partitions.
The moduli space of quasimaps to $\mathcal{M}(\mathsf{v},\mathsf{w})$ which are fixed by $\mathsf{A}$ is isomorphic to the moduli space of quasimaps to $\mathcal{M}(\mathsf{v},\mathsf{w})^{\mathsf{A}}$, see [46] section 7.3. By Proposition 3.2, it suffices to describe torus fixed quasimaps when $|\mathsf{w}|=1$. In this case, the fixed point $x$ is given by a single partition $\lambda$.
As recalled above, a quasimap from $\mathbb{P}^{1}$ to $\mathcal{M}(\mathsf{v},\mathsf{w})$ provides the data of vector bundles $\mathscr{V}_{i}$. These split into a direct sum of line bundles $\mathscr{V}_{i}=\bigoplus_{j}\mathcal{O}(d_{i,j})$ where $0\leq i\leq l-1$ and $1\leq j\leq\mathsf{v}_{i}$. If the quasimap is $\mathsf{T}\times\mathbb{C}^{\times}_{q}$-fixed, then this splitting is additionally graded by the $\mathsf{T}$-weights of $\mathcal{V}_{i}|_{\lambda}$ which thus puts the numbers $d_{i,j}$ into canonical bijection with boxes $\square\in\lambda$. One can additionally show, see for example [12] section 5.1.3, that the integers $d_{i,j}$ must form a “reverse plane partition” over $\lambda$. In our notation, this means that we have a collection of non-negative integers $d_{\square}$ for each $\square\in\lambda$ such that $d_{\square}\leq d_{\square^{\prime}}$ whenever $\square^{\prime}$ is immediately above or right of $\square$, with directions as in section 2.1. Furthermore, the reverse plane partition uniquely determines the $\mathsf{T}\times\mathbb{C}^{\times}_{q}$-fixed quasimap.